# Residual Diffusion Modeling for Km-scale Atmospheric Downscaling

### v3

@@ 3.3 Case studies: downscaling coherent structures
Operational meteorologists value case study analysis, since aggregate skill scores and spectra can be more easily gamed or mask symptoms of spatial incoherence. We thus turn our attention to specific weather regimes. Fig. 3 illustrates the variability of the generated radar reflectivity field at four distinct times. When analyzing 200 samples, the standard deviation of radar reflectivity (second column from the left) is roughly 20% of the magnitude of the mean (left column), with the majority of the variance located in and around existing precipitation regions. Such a pattern is anticipated given that the timing and location of precipitation can vary markedly even within a fixed large-scale configuration since convective physics are inherently stochastic.
The CorrDiff prediction for an individual sample (number 200, third column from the left) reveals a fine-scale structure akin to the target data (right column). The similarity between panels (a) and (d) highlights the role of the mean prediction in forming large-scale coherent structures, such as typhoon Chanthu (2021), top row, and frontal systems, bottom row. Specifically the typhoon rainbands (spiral bands of clouds that emanate from the typhoon center) are coherent enough to be captured by the mean, but with large variance and smooth structure. The fine-scale structure reflecting the stochastic physics is captured well by the diffusion model, refining the smooth fields of the mean prediction as seen in the third column of Fig. 3. Further comparison across samples is presented in an animation in S3 in 4.
Figure 3: Demonstration of the stochastic prediction of radar reflectivity (in dBZ). Top to bottom: 2021-09-12 00:00, 2021-04-02 06:00, 2021-02-02 12:00 and 2022-02-13 20:00. Left to right: sample mean, sample standard deviation, sample number 200 and the target forecast.
@@ 3.3.1 Frontal system case study
Frontal systems are an example of organized atmospheric systems. A cold front is a sharp change in temperature and winds associated with a mature cyclonic storm. As the front moves eastward, the cold air pushes the warm air to its east upward. This upward motion leads to cooling, condensation and ultimately rainfall. That is, these physics should manifest as multi-variate relationships with linked fine scale structures of two wind vector components and temperature that should co-locate with radar reflectivity.
Fig. 4 shows an example of CorrDiff downscaling a cold front. The position of the front is clearly visible in the southeast portion of the domain, where a strong horizontal surface temperature gradient (top) co-locates with strong across-front wind convergence (middle). The along-front components of the wind vector also change abruptly (middle row), which is consistent with the change in temperature. The super resolved gradients in the temperature and winds are encouragingly sharper than the input. The intense rainfall associated with this convergence line can be seen in the radar reflectivity ground truth for the same calendar date (which are shown in bottom row of Fig. 3 above). The generated radar reflectivity is appropriately concentrated near the frontal boundary. Both the location of the front, and the magnitude of the horizontal wind and temperature gradients (sharpening of the front) associated with it, are captured well by CorrDiff although some mispositioning of the exact front location is inevitable. These are reassuring signs of learnt physics during the generation task.
Figure 4: Examining the downscaling of a cold front on Feb 2, 2022 at 20 UTC. Left to right: prediction of ERA5, CorrDiff and Target for different fields, followed by their averaged cross section from 20 lines parallel to the thin dashed line in the contour figures. Top to bottom: 2 meter temperature (arrows are true wind vectors), along front wind (arrows are along front wind component) and across front wind (arrows are across front wind component).
@@ 3.3.2 Tropical Cyclone case study
Downscaling typhoons (i.e. tropical cyclones) is especially complicated. The average radius of maximum winds of a tropical cyclone is less than 100km, and at 25km resolution of the input data tropical cyclones are only partially resolved, resulting in cyclonic structures that are too wide and too weak compared with high resolution models or observations [8]. A useful downscaling model must simultaneously correct their size and intensity in addition to generating appropriate fine-scale structure.
We find that CorrDiff is able to correct the structure of tropical cyclones accurately. Fig. 5(a)-(f) shows the structure of typhoon Chanthu (2021) on September 12 at 00:00:000 UTC. Compared to the target data (panel c) the poorly resolved typhoon in the ERA5 (panel a) is too wide and does not include a closed contour annulus of winds above 16 m/s surrounding an overly quiescent eye-wall. CorrDiff downscaling (panel b) is able to recover much of the spatial structure of the windspeed compared with the target. The skill of the CorrDiff downscaling compared to interpolating ERA5 can be more clearly quantified by calculating the mean axisymmetric structure of the storms as a function of radius from eye-wall center (panel f). Notably, with downscaling the radius of maximum winds decreases from 75km to about 25km while the windspeed increases from 20 m/s to 50 m/s – both favorable improvements.
@@
Figure 5: A comparison of the 10m windspeed maps (m/s), distributions and the axisymmetric cross section from typhoon Chanthu (2021) on 2021/09/11:12:00:00UTC. Panels (a),(b),(c) show the 10m windspeed from ERA5, CorrDiff downscaling of ERA5 and the target (WRF), respectively. The CorrDiff panels show the first ensemble member. The solid black contour indicates the Taiwan coastline. Storm center of the ERA5, CorrDiff and WRF are shown in red ‘+‘, orange diamond, and the black dot, respectively. Panels (d) and (e) show the distribution shift (normalized PDFs) for the entire CWA domain and for the typhoon selected region in the top panels. Panel (f) shows the axisymmteric structure of the typhoon about its center. For the CorrDiff curves, line is the ensemble mean and the shading shows one standard deviation around the mean.
Probabilities of damaging typhoon winds shown in in panels (d) and (e) in Fig. 5 are significantly improved. In ERA5 (red), occurrence of weak wind speeds is over-estimated and the damaging extreme hurricane winds (above 25m/s) are missing. CorrDiff better predicts the chances of the strong winds most likely to impact society and infrastructure. Further analysis 5 expands beyond this case study to examine generated wind statistics for several hundred typhoons that crossed the domain during 1980 to 2020. Although no target data is available, when comparing the maximum windspeed and radius of maximum windspeed from CorrDiff predictions to a reference from the Japan Meteorological Agency best track dataset [4] CorrDiff is found to correct at least 75% of the error in intensity for windspeed below 50m/s but only 50% of the error for higher windspeeds S4.
3.4 Downscaling a forecast from a global model
An attractive use case for CorrDiff is to directly downscale forecasts from global models. During the writing of this paper, typhoon Koinu (2023) approached Taiwan. Together with CWA we compared the live forecasts of Koinu from GFS, CorrDiff downscaling applied to GFS (here we use the same input channel as before) and WRF (target) data, see Figure 6. Expectation of good results should be tempered by the fact that this exposes CorrDiff to out of sample inputs, by using GFS in place of ERA5. Even without retraining to include the leadtime forecast bias of GFS, we find some encouraging results. An expected enhancement in windspeed intensity is confirmed and it most accurate up to a leadtime of 6 hours (third row) after which is decreases. The radar reflectivity shows some caveats. While the domain averaged reflectivity is corrected well by CorrDiff, the coherent rainband that is exhibited in the target data at lead times larger than 3h is absent in CorrDiff. This rainband is also absent in the input (GFS), and CorrDiff seem to struggle to generate it. This suggests that CorrDiff could be used for downscaling from a global model at short lead times (12h, for which baseline was provided).
Figure 6: A comparison of a live km-scale forecast made by linking GFS predictions to CorrDiff donwscaling, validated against WRF-CWA operational predictions made prior to the typhoon Loinu landfall. Forecast is initialized on 20231004-12:00 UTC, top to bottom show lead time 0, 3, 6, 9, 12 hours. Left to right: forecast from the GFS; CorrDiff downscaling of GFS and the target (WRF-CWA). The fourth column compares the axisymmetric windspeed profile of the typhoon (shading for GFS+CorrDiff shows one standard deviation). The domain mean of the 1-hr maximum radar reflectivity is indicated at the top left of each of the reflectivity maps.
@@ 4 Discussion
This study presents a generative diffusion model (CorrDiff) for downscaling coarse-resolution (25-km) global weather state to higher resolution (2km) over a subset of the globe. CorrDiff consists of two steps: regression and generation. The regression step approximates the mean, while the generation step further corrects the mean but also generates the distribution, producing fine-scale details stochastically. This approach is akin to the decomposition of physical variables into their mean and perturbations, common practice in fluid dynamics, e.g. [40].
Through extensive testing in the region of Taiwan, the model is shown to skillfully correct kinetic energy spectra, generate realistic probabilities and downscale coherent structures accurately(with minor caveats such as inaccuracies in the 2-meter temperature on the warm sector of a frontal system (Fig. 4) and over-correction of typhoon sizes in some out-of-sample cases for which we do not have the target (WRF) data, see 5). It is possible that the model’s accuracy could be further improved with a larger training dataset that contains more diverse examples of such rare coherent structures.
@@
The two step approach in CorrDiff also offers the possibility to trade off between the fast inference of the mean using the UNet-regression, and the accurate and probabilistic inference of the CorrDiff. This is particularly useful given that some variables – like the 2 meter temperature – are well predicted by the UNet-regression while others like radar reflectivity depend on the diffusion step for their skill (see Figure 2). Moreover, it could be possible to apply the diffusion step to a mean prediction obtained in a different way (e.g. a numerical model if available) to generate a plausible distribution from a single prediction.
With the current hardware (see 6) and with an unoptimized code (no GPU parallelization or batching utilized), CorrDiff inference is about $22$ ($40$) times faster, and $1,130$ ($1,800$) times more energy efficient than running CWA-WRF on CPU with float32 (mixed) precision.
@@
This paper focused on generation quality, and not on optimal inference speed, for which gains could be easily anticipated. Our prototype of CorrDiff is using a dozen iterations thanks to the initial regression step. However, for future work, we will push to reduce the number of iterations to only a few by using distillation methods [49, 61, 62] and pursuing other performance optimization techniques [34, 58].
@@
Several potential extensions of the proposed method are worth considering:
@@
    Downscaling Coarse-Resolution Medium-Range Forecasts: This requires addressing lead time-dependent forecast errors, enabling a comprehensive evaluation of simultaneous bias correction and downscaling.
@@ 5.1 Background on diffusion models
SDE formulation. To present the forward and backward processes rigorously, they can be captured via stochastic differential equations (SDEs). Such SDEs ensure that the sample, $\bf x$, aligns with the designated data distribution, $p$, over its progression through time [55, 28]. A numerical SDE solver can be used here, where a critical component is the noise schedule, $\sigma(t)$, which prescribes the noise level at a specific time, $t$. A typical noise scheduler is $\sigma(t)\propto\sqrt{t}$. Based on [28], the forward SDE is given as
@@
The term $\dot{\sigma}(t)$ refers to the time derivative of $\sigma(t)$. The forward SDE is a Wiener process, while the backward SDE comprises two terms: a deterministic component representing the probability flow ODE with noise degradation, and noise injection via the Wiener process.
@@ 5.2 Proposed approach
As discussed in section 2, the high resolution state ${\bf{x}}$ in (1) can be written as the sum of the mean $\boldsymbol{\mu}$ and the difference $\bf{r}$, where the latter will be nearly zero mean and exhibits a small distribution shift, which facilities training diffusion models for correcting the mean prediction. It is worth noting that this two-step method has further implications for learning physics. The UNet-regression step can anticipate many of the physics of downscaling, some of which are deterministic. These include high-resolution topography (which to first order controls the 2-meter temperature variation due to the lapse-rate effect), and the large-scale horizontal wind which combine leading balances in the free atmosphere with the effect of surface friction and topography. Stochastic phenomena such as convective storms that also change temperatures and winds are easier to model as deviations from the mean. Also, cloud resolving models are explicitly formulated using deviations from a larger scale balanced state [42]. In the next section, we discuss the regression and generation step in details.

### v4

@@ 3.5 Case studies: downscaling coherent structures
We now turn our attention to specific weather regimes, which are important to examine since aggregate skill scores and spectra can be more easily gamed and mask symptoms of spatial incoherence. Fig. 4 illustrates the variability of the generated radar reflectivity field on four separate dates corresponding to distinct Taiwanese weather events. Two dates are chosen randomly (Fig. 4 e-k). The other two correspond to a dates where coherent events such as typhoon (Fig. 4 a-d) and frontal event (Fig. 4 m-o) are present; these dates are further analyzed in the following sections and more examples of both of these phenomena are provided in the Appendix for additional context 6. The standard deviation across our ensemble of 32 generated CorrDiff samples (second column from the left) is roughly 20% of the magnitude of the ensemble mean (left column). The CorrDiff prediction for an arbitrary ensemble member (last sample; 32nd member; third column from the left) is useful to compare to the target data (right column). However, due to the stochastic nature of the generation, some disagreement in detailed patterns and positioning should be expected. The similarity between the first and the third columns highlights the role of the mean UNet prediction in forming large-scale coherent structures, such as the positioning of rainbands within typhoon Haikui (2023), top row, and frontal systems, bottom row. The additional fine-scale structure reflecting the stochastic physics contributed by the diffusion model is seen in the third column of Fig. 4. Further comparison across independent generated samples is presented in an animation in S4 in 4 that is helpful for appreciating the portion of the generated image that is governed by the corrective diffusion subcomponent of CorrDiff.
Figure 4: Demonstration of the stochastic prediction of radar reflectivity (in dBZ). Top to bottom: 2023-09-03 00:00:00 , 2021-02-17 21:00:00, 2021-03-04 01:00:00 and 2022-02-13 20:00:00 UTC. Left to right: sample mean, sample standard deviation, sample number 32 and the target forecast.
@@ 3.5.1 Frontal system case study
Frontal systems are an example of organized atmospheric systems. A cold front is a sharp change in temperature and winds associated with a mature, mid-latitude, cyclonic storm. As the front moves eastward, the cold air pushes the warm air to its east upward. This upward motion leads to cooling, condensation and ultimately rainfall. That is, these physics should manifest as multi-variate relationships with linked fine scale structures of two wind vector components and temperature that should co-locate with radar reflectivity.
Fig. 5 shows an example of CorrDiff downscaling a cold front. Examining the target data (WRF in third column), the position of the front is clearly visible in the southeast portion of the domain, where a strong horizontal 2-meter temperature gradient (top) co-locates with both a strong divergence of the across-front wind (bottom) and a reversal in direction of the along-front wind on either side of the temperature front (middle). Compared to the target data the ERA5 representation of this front is smoother. CorrDiff partially restores sharpness to the front by increasing the horizontal gradients across all three field variables. Although the generated front has some differences in morphology compared to the ground truth, the consistency of its morphology across winds and temperature is reassuring. The intense rainfall associated with the convergence at the front can be seen in the radar reflectivity for the same date and time in bottom row of Fig. 4. The generated radar reflectivity is appropriately concentrated near the sharpened frontal boundary at the cold sector. We expand this analysis of the frontal boundaries across more samples in Section 6.1, which reveals that CorrDiff consistently adjusts the winds, temperature and radar reflectivity at the front, but that its skill in sharpening frontal gradients exhibits case-to-case variability.
Figure 5: Examining the downscaling of a cold front on 2022-02-13 20:00:00 UTC. Left to right: prediction of ERA5, CorrDiff and Target for different fields, followed by their averaged cross section from 21 lines parallel to the thin dashed line in the contour figures. Top to bottom: 2-meter temperature (arrows are wind vectors), along front wind (arrows are along-front component of the wind vector) and across front wind (arrows are across-front component of the wind vector). At the right column the cross sections of the WRF (black line) and ERA5 (red line) are compared with the mean of a 32 member ensemble prediction from CorrDiff (orange line) where the shading shows $\pm$ one standard deviation.
@@ 3.5.2 Tropical Cyclone case study
Downscaling typhoons (i.e., tropical cyclones) is especially complicated, helpfully revealing the limitations of CorrDiff for representing extreme events. Not only are typhoons extremely rare in our training data, but the average radius of maximum winds of a tropical cyclone is less than 100km such that at the 25km resolution of our input data tropical cyclones are only partially resolved. This leads to a cyclonic structures that are too wide in horizontal extent and too weak in wind intensity compared with high-resolution models or observations [8]. A useful downscaling model should simultaneously correct their size and intensity in addition to generating appropriate fine-scale structure.
An example illustrating the benefits and limitations of CorrDiff for downscaling typhoon Haikui (2023) is shown in Figure 6. Compared to the ground truth (Fig. 6d), the ERA5 reanalysis (Fig. 6a) poorly resolves the typhoon, depicting it as overly wide and with no closed contour annulus of winds above $16ms^{-1}$. The UNet (Fig. 6b) likewise fails to recover a closed contour, although it does helpfully corrects approximately 50% of error in the large-scale windspeed and structure compared to the target. CorrDiff (Fig. 6c) enhances the UNet by adding spatial variability, but maintains similar intensity.
The benefits of the CorrDiff downscaling compared to interpolating ERA5 can be more clearly quantified by examining the logarithm of the PDF of the windspeed, Fig. 6(e). In the ERA5 the wind speed PDF has a sharp cutoff such that high wind speed values in excess of $27ms^{-1}$ are missing. CorrDiff partially restores the tail of the typhoon wind speed PDF, and is capable of predicting wind speeds up to $40ms^{-1}$ compared with the maximum value of $50ms^{-1}$ in the target. The diffusion component of CorrDiff is responsible for the most extreme wind speeds in its predictions. In contrast, the mean axisymmetric structure of the typhoon (Fig. 6f), is controlled more by the UNet, which reveals the influence of CorrDiff on typhoon geometry: With downscaling the radius of maximum winds decreases from 75km in ERA5 to about 50km in CorrDiff, compared with 25km in the WRF model. At the same time, the axisymmetric windspeed maximum increases from $22ms^{-1}$ in ERA5 to $33ms^{-1}$, compared with 45 in WRF – both favorable improvements. Ultimately, CorrDiff is able to synthesis consistent radar reflectivity (see top row of Fig. 4.
@@
Figure 6: A comparison of the 10m windspeed maps ($ms^{-1}$), distributions and the axisymmetric cross section from typhoon Haikui (2023) on 2023-09-03 00:00:00 UTC. Panels (a),(b),(c), (d) show the 10m windspeed from ERA5, UNet, CorrDiff and target (WRF), respectively. The CorrDiff panels show the first of 32 ensemble members. The solid black contour indicates the Taiwan coastline. Storm center of the ERA5, CorrDiff and WRF are shown in red ‘+‘, orange diamond, and the black dot, respectively. Panels (e) shows the logarithm of the PDF of windspeed. Panel (f) shows the axisymmteric structure of the typhoon about its center, where for the CorrDiff curves, solid line shows the ensemble mean and the shading shows $\pm$ one standard deviation along the ensemble dimension.
For further discussion of typhoon downscaling, we refer the interested reader to Appendix Section 6.2, where we explore additional date-times for Haikui (2023), investigate an additional typhoon Chanthu (2021), and analyze generated wind statistics across a 600-member ensemble of typhoon-containing time intervals spanning the 1980-2020 period. This extended analysis suggests that, while the main results emphasized for are case study above are qualitatively representative when typhoons are far from land, the diffusion component of CorrDiff frequently plays a stronger role in intensifying typhoon axisymmetric structure, and CorrDiff tends to lead to too much horizontal contraction of cyclone morphology, predicting a radius of maximum winds that is statistically too small.
@@ 4 Discussion
This study presents a generative diffusion model (CorrDiff) for multivariate downscaling of coarse-resolution (25-km) global weather states to higher resolution (2km) over a subset of the globe, and simultaneous radar channel synthesis. CorrDiff consists of two steps: regression and generation. The regression step approximates the mean, while the generation step further corrects the mean but also generates the distribution, producing fine-scale details stochastically. This approach is akin to the decomposition of physical variables into their mean and perturbations, common practice in fluid dynamics, e.g. [47].
Through extensive testing in the region of Taiwan, the model is shown to produce reasonably realistic power spectra and probability distributions of all target variables. The diffusion component of CorrDiff is found to be especially important for the task of radar channel synthesis. Several case studies reveal that the model is able to downscale coherent structures consistently across its variables. Focusing on a midlatitude frontal event, horizontally co-located gradients of winds and temperatures are generated alongside spatially consistent radar reflectivity features, with incomplete but improved sharpness. For typhoons, encouraging partial corrections of typhoon size and wind speed intensity are found, alongside generated radar echos containing qualitatively realistic km-scale details reminiscent of tropical cyclone rainband morphology. It is logical to expect that the model’s accuracy could be further improved with a larger training dataset that contains more diverse examples of such rare coherent structures such as by pre-training on large libraries of typhoons generated by high-resolution physical simulators; we encourage work in this direction.
Another important unsolved challenge is optimally calibrating CorrDiff’s generated uncertainty to better match its error levels. This is somewhat unexpected since diffusion models for image generation are known to be over-dispersive in the sense of producing low-quality samples and variance-reducing techniques are often used during the sampling to discourage such outliers [30, 34]. The lack of grid-point-level spread here could owe to a number of factors in the diffusion training process including the noise schedules used, the comparatively large resolution (448x488) compared to typical image generation (64x64), or the weighting in the loss function.
To become useful for km-scale weather prediction, extensions of CorrDiff are encouraged that include temporal coherence, such as via video diffusion or learnt autoregressive km-scale dynamics; as with super-resolution these must be formulated as stochastic machine learning tasks. Currently, beyond the coherence of the large scale conditioning given from ERA5 there is no guarantee that CorrDiff’s km-scale dynamics will be coherent in time. Additional integrations with km-scale data assimilation are also essential for this use case. Our current demonstration relies only on the global data assimilation (DA) used to produce the ERA5 dataset. Unlike the target data it is trained on, CorrDiff effectively bypasses the regional DA, which for CWA is of similar computational cost as running the operational numerical model (WRF) model for 13h.
@@
For such extensions, the two step approach in CorrDiff offers practical advantages to reduce the amount of variance that must be handled stochastically, and trade-off between the fast inference of the mean using the UNet, and the probabilistic inference of the CorrDiff. This is particularly useful given that some variables depend more than others on the diffusion step for their skill (see Figure 2). Moreover, it could be possible to apply the diffusion step to a mean prediction obtained in a different way (e.g. a numerical model if available) to generate a plausible distribution from a single prediction.
@@
With the current hardware and code-base CorrDiff inference is about $652$ times faster, and $1,310$ times more energy efficient than running CWA-WRF on CPUs, although such a comparison between dynamical and statistical downscaling is limited (see 6.3 for details). This paper focused on generation quality, and not on optimal inference speed, for which gains could be easily anticipated. Our CorrDiff prototype is using a dozen iterations thanks to the initial regression step. Refinement of the technique could reduce the number of iterations to only a few by using distillation methods [59, 73, 74] and pursuing other performance optimization techniques [41, 69].
@@
If some of the above challenges in the model are resolved, several potential extensions of the proposed method are worth consideration by the community:
@@
    Downscaling Coarse-Resolution Medium-Range Forecasts: This requires addressing lead time-dependent forecast errors in the input, enabling a comprehensive evaluation of simultaneous bias correction and downscaling, and adding temporal coherence and km-scale prediction and data assimilation capabilities.
@@
  4. 4.
    Synthesizing sub-km sensor observations: To achieve effective resolutions beyond what is possible to simulate today, and sidestep issues of numerical simulation, it would be interesting to test whether variants of CorrDiff can be trained to generate raw senor observations where dense networks exists. Our demonstrated ability to synthesize an observable as challenging as radar reflectivity from column water vapor should embolden such efforts.
@@ 5.1 Background on diffusion models
formulation of the underlying stochastic differential equations. To present the forward and backward processes rigorously, they can be captured via stochastic differential equations (SDEs). Such SDEs ensure that the sample, $\bf x$, aligns with the designated data distribution, $p$, over its progression through time [65, 33]. A numerical SDE solver can be used here, where a critical component is the noise schedule, $\sigma(t)$, which prescribes the noise level at a specific diffusion-time, $t$. Here ’diffusion time’ is a virtual variable used for denoting denoising steps and has roots in differential equations that are derived from Langevin dynamics. To avoid confusion with the time of day, we use diffusion time to denote this variable. A typical noise scheduler is $\sigma(t)\propto\sqrt{t}$. Based on [33], the forward SDE is given as
@@
The term $\dot{\sigma}(t)$ refers to the derivative of $\sigma(t)$ with respect to the diffusion-time. Here $\omega$ in the forward SDE is a Wiener process, while the backward SDE comprises two terms: a deterministic component representing the probability flow ODE with noise degradation, and noise injection via the Wiener process.
@@ 5.2 Proposed approach
As discussed in section 2, the Fig. state ${\bf{x}}$ in (1) can be written as the sum of the mean $\boldsymbol{\mu}$ and the difference $\bf{r}$, where the latter will be nearly zero mean and exhibits a small distribution shift, which facilities training diffusion models for correcting the mean prediction. It is worth noting that this two-step method has further implications for learning physics. The UNet-regression step can anticipate many of the physics of downscaling, some of which are deterministic. These include high-resolution topography (which to first order controls the 2-meter temperature variation due to the lapse-rate effect), and the large-scale horizontal wind which combine leading balances in the free atmosphere with the effect of surface friction and topography. Stochastic phenomena such as convective storms that also change temperatures and winds are easier to model as deviations from the mean. Also, cloud resolving models are explicitly formulated using deviations from a larger scale balanced state [49]. In the next section, we discuss the regression and generation step in details.